OpenStudy (anonymous):

Will give medal and fan Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses.

OpenStudy (anonymous):

f(x)= x+a/b g(x)=cx+d

OpenStudy (anonymous):

@agent0smith please help.

OpenStudy (agent0smith):

Start with f(x), pick numbers for a and b, like 1 and 2 and replace a and b with with them. Then find the inverse of f(x), since that'll help you get c and d.